# Predictable Scale: Part I — Optimal Hyperparameter Scaling Law in Large Language Model Pretraining

### v1

@@ Abstract
The impressive capabilities of Large Language Models (LLMs) across diverse tasks are now well-established, yet their effective deployment necessitates careful hyperparameter optimization. Through extensive empirical studies involving grid search across diverse configurations, we discover universal scaling laws governing these hyperparameters: optimal learning rate follows a power-law relationship with both model parameters and data sizes, while optimal batch size scales primarily with data sizes. Our analysis reveals a convex optimization landscape for hyperparameters under fixed models and data size conditions. This convexity implies an optimal hyperparameter plateau. We contribute a universal, plug-and-play optimal hyperparameter tool for the community. Its estimated values on the test set are merely 0.07% away from the globally optimal LLM performance found via exhaustive search. These laws demonstrate remarkable robustness across variations in model sparsity, training data distribution, and model shape. To our best known, this is the first work that unifies different model shapes and structures, such as Mixture-of-Experts models and dense transformers, as well as to establish optimal hyperparameter scaling laws across diverse data distributions. This exhaustive optimization process demands substantial computational resources, utilizing nearly one million NVIDIA H800 GPU hours to train 3,700 LLMs of varying sizes and hyperparameters from scratch and consuming approximately 100 trillion tokens in total. To facilitate reproducibility and further research, we will progressively release all loss measurements and model checkpoints through our designated repository https://step-law.github.io/ .
@@ 1 Introduction
State-of-the-art Large Language Models (LLMs) Brown et al. (2020); Jin et al. (2023); Biderman et al. (2023); Scao et al. (2022); Touvron et al. (2023a, b); Grattafiori et al. (2024a); DeepSeek-AI et al. (2024a); Yang et al. (2024); DeepSeek-AI et al. (2024b, 2025), have reached unprecedented scales, with models being trained on billions of parameters and trillions of tokens. Recent developments like Llama 3 Grattafiori et al. (2024a) demonstrate this trend, utilizing 15 trillion tokens for training Grattafiori et al. (2024a). At such massive scales, identifying optimal hyperparameter configurations becomes both critical for model performance and challenging due to computational constraints.
@@ 2 Related Works
In theory-driven approaches, $\mu\text{P}$ Yang et al. (2022) established foundational learning rate transfer rules for varying model widths, though this required specific modifications to initialization and attention mechanisms. However, the $\mu\text{P}$ framework and its extensions Yang et al. (2022); Everett et al. (2024); Lingle (2024); Blake et al. (2024); Yang et al. (2023); Bordelon et al. (2023) are limited in scope, lacking guidance for learning rate adaptation across different data distributions, model architectures, sparsity levels, and data sizes. Additionally, these works do not address batch size optimization.
@@ 4.1.1 Evaluation metric
As shown in Fig. 3, for the case where the smoothed training loss converges to the optimal value of $2.279$ ( as indicated by the solid red-framed line in Fig. 3 (b)), the corresponding LR and BS are $1.95\times 10^{-3}$ and $393,216$ respectively. This is the same as the position of the LR and BS corresponding to the validation loss converging to the optimal value of $2.038$ (as indicated by the solid red-framed line in Fig. 3 left). Moreover, the overall trend of how the smoothed training loss deviates from the optimal value with varying learning rates and batch sizes (as shown by the heatmap patterns in Fig. 3 right) closely mirrors the corresponding variations observed in validation loss measurements. This alignment demonstrates that the smoothed training loss provides consistent optimization guidance for learning rate and batch size selection, matching the parameter configurations that would be obtained through direct validation loss evaluation.
@@ 5.2 Sparsity-Independent in MoE
Figure 8: Validation loss landscapes of MoE models under varying sparsity ratios ($N_{a}/N$). Left: Low sparsity ($N_{a}/N=0.27$). Middle: Medium sparsity ($N_{a}/N=0.58$). Right: Reduced model depth ($D=8.0B$) at medium sparsity. Our method consistently approximates global minima across sparsity regimes.
@@ A.1 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
A.2 Model Structural Parameters
@@
Table 5: MoE Model Configuration.
@@
Table 6: Comparison of dataset weights (%) across different training recipes. Each recipe represents a different focus: baseline (llama1), enhanced code and mathematics capability, and English-Chinese bilingual ability.
A.3 Dense Models Results

### v4

@@ Abstract
The impressive capabilities of Large Language Models (LLMs) across diverse tasks are now well-established, yet their effective deployment necessitates careful hyperparameter optimization. Through extensive empirical studies involving grid searches across diverse configurations, we discover universal scaling laws governing these hyperparameters: optimal learning rate follows a power-law relationship with both model parameters and data sizes, while optimal batch size scales primarily with data sizes. Our analysis reveals a convex optimization landscape for hyperparameters under fixed models and data size conditions. This convexity implies an optimal hyperparameter plateau. We contribute a universal, plug-and-play optimal hyperparameter tool for the community. Its estimated values on the test set are merely 0.09% away from the globally optimal LLM performance found via an exhaustive search. These laws demonstrate remarkable robustness across variations in model sparsity, training data distribution, and model shape. To our best known, this is the first work that unifies different model shapes and structures, such as Mixture-of-Experts models and dense transformers, as well as establishes optimal hyperparameter scaling laws across diverse data distributions. This exhaustive optimization process demands substantial computational resources, utilizing nearly one million NVIDIA H800 GPU hours to train 3,700 LLMs of varying sizes and hyperparameters from scratch and consuming approximately 100 trillion tokens in total. To facilitate reproducibility and further research, we will progressively release all loss measurements and model checkpoints through our designated repository https://step-law.github.io/.
@@ 1 Introduction
State-of-the-art Large Language Models (LLMs) Brown et al. (2020); Jin et al. (2023); Biderman et al. (2023); Scao et al. (2022); Touvron et al. (2023a, b); Grattafiori et al. (2024); DeepSeek-AI et al. (2024a); Yang et al. (2024); DeepSeek-AI et al. (2024b, 2025), have reached unprecedented scales, with models being trained on billions of parameters and trillions of tokens. Recent developments like Llama 3 Grattafiori et al. (2024) demonstrate this trend, utilizing 15 trillion tokens for training Grattafiori et al. (2024). At such massive scales, identifying optimal hyperparameter configurations becomes both critical for model performance and challenging due to computational constraints.
@@ 2 Related Works
In theory-driven approaches, $\mu\text{P}$ Yang et al. (2022) established foundational learning rate transfer rules for varying model widths, though this required specific modifications to initialization and attention mechanisms. However, the $\mu\text{P}$ framework and its extensions Yang et al. (2022); Everett et al. (2024); Lingle (2024); Blake et al. (2024); Yang et al. (2023); Bordelon et al. (2023) are limited in scope, lacking guidance for learning rate adaptation across different data distributions, model architectures, sparsity levels, and data sizes. Additionally, these works do not address optimal batch size prediction.
@@ 4.1.1 Evaluation metric
As shown in Fig. 3, for the case where the smoothed training loss converges to the optimal value of $2.279$ ( as indicated by the solid red-framed line in Fig. 3 right) , the corresponding LR and BS are $1.95\times 10^{-3}$ and $393,216$ respectively. This is the same as the position of the LR and BS corresponding to the validation loss converging to the optimal value of $2.038$ (as indicated by the solid red-framed line in Fig. 3 left). Moreover, the overall trend of how the smoothed training loss deviates from the optimal value with varying learning rates and batch sizes (as shown by the heatmap patterns in Fig. 3 right) closely mirrors the corresponding variations observed in validation loss measurements. This alignment demonstrates that the smoothed training loss provides consistent optimization guidance for learning rate and batch size selection, matching the parameter configurations that would be obtained through direct validation loss evaluation.
@@ 5.2 Sparsity-Independent in MoE
Figure 8: Validation loss landscapes of MoE models under varying sparsity ratios ($N_{a}/N$). Left: Low sparsity ($N_{a}/N=0.27$). Middle: Medium sparsity ($N_{a}/N=0.58$). Right: Medium sparsity at D=8.0B. Our method consistently approximates global minima across sparsity regimes.
@@ A.1 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
A.2 Model Structural Parameters And Results
@@
Table 5: MoE Model Configuration. $N_{{expert}}$ denotes the number of experts. $d_{{moe}}$ denotes the hidden size of experts. Top-$k$ denotes the number in the routing algorithm. $N_{a}$ denotes the activate parameters.
@@
Table 6: Comparison of dataset weights (%) across different training recipes. Each recipe represents a different focus: baseline, enhanced code and mathematics capability, and English-Chinese bilingual ability.
@@
Figure 41: Illustration of Hyperparameter Configuration Space for MoE Model 12.
Figure 42: Illustration of Hyperparameter Configuration Space for MoE Model 13.
Figure 43: Illustration of Hyperparameter Configuration Space for MoE Model 14.
Figure 44: Illustration of Hyperparameter Configuration Space for MoE Model 15.